# A Systematic Literature Review of Parameter-Efficient Fine-Tuning for Large Code Models

### v2

@@ 4.2.2. PEFT Methods Across Software Engineering Tasks and Model Architectures
In this section, we explore how PEFT methods are distributed across various software engineering tasks, distinguishing them between generative and non-generative. Among the identified techniques, Base LoRA is the most dominant, appearing in 24 SE tasks. The Adapter family has been applied across 23 software engineering tasks, encompassing specialized variants such as L-Adapters and T-Adapters. Other frequently used techniques include Prompt Tuning (14 studies), Prefix Tuning (13), and P-Tuning (7), which leverage soft or continuous prompts to inject task-specific knowledge into frozen pre-trained models. Collectively, these methods demonstrate significant effectiveness in reducing training costs while maintaining or enhancing performance, affirming their value in scalable and efficient model adaptation across a diverse set of software engineering tasks.
@@
Code Generation is addressed in 9 studies, making it one of the most frequently explored generative SE tasks in the context of PEFT. LoRA remains the dominant technique, appearing in 6 studies (Wang et al., 2025; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Ayupov and Chirkova, 2022; Yang et al., 2024a), including combinations of FF-LoRA variants with Adapter Tuning, one study implemented Code Summarization, Code Clone Detection, Code Generation, and Code Translation (Ayupov and Chirkova, 2022).
@@
A further application of SE automation, revolves around the generation of unit tests–referred with Unit Test Generation in our taxonomy Fig. 3. In particular, for such a task we have found two distinct studies, each employing distinct PEFT approaches. Liu et al. (Liu et al., 2024a) applied QLoRA within a multi-task fine-tuning framework (MFT-5Tasks), while Zlotchevski et al. (Zlotchevski et al., 2022) employed Prefix Tuning, wherein 200 trainable virtual tokens were prepended to a frozen BART model.
@@
Looking at SE tasks for which code changes optimization are requested, we find a surprisingly widespread application of PEFT techniques to support Code Refinement activities (i.e., bug-fixing). This peculiar task, has been explored through a diverse range of PEFT techniques, which—consistent with findings from Code Generation—span the full spectrum of parameter-efficient optimization strategies. In this context, Nuo et al. (Chen et al., 2023b) introduced Pass-Tuning, a lightweight method that tunes less than 1% of parameters by leveraging U-AST (Upgraded Abstract Syntax Tree) token sequences within a graph-attentional model. Building on structure-aware prompting, Yun et al. (Peng et al., 2024) applied Prompt Tuning to CodeT5-base, crafting domain-specific prompts based on AST-derived fix templates. Saberi et al. (Saberi and Fard, 2023) proposed a fusion-based approach, combining a Java-specific L-Adapter with a token-type-aware NER-Adapter via AdapterFusion to improve bug-fixing effectiveness. Similarly, Hajipour et al. (Hajipour et al., 2024) investigated the use of LoRA as a parameter-efficient alternative for Code Refinement tasks. Traditional Prompt Tuning appears in one study (Peng et al., 2024), while a multi-perspective adaptation strategy—incorporating L-Adapter, NER-Adapter, and AdapterFusion—is explored in a different research (Saberi and Fard, 2023). Additionally, one study applies a comprehensive suite of PEFT methods—including Pass-Tuning, BitFit, Adapters, and P-Tuning—to enhance source code refinement (Chen et al., 2023b).
@@
Still in the domain of Code2Code tasks, Code Translation is addressed in 7 studies (Choi and Lee, 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b; Liu et al., 2024a), while Code Completion is explored in 2 studies (Shi et al., 2023; Liu et al., 2024a). Notably, for the latter, Shi et al. (Shi et al., 2023) introduce a customized PEFT technique called Telly-K, which modifies the training procedure by selectively freezing the lower layers of the model. This approach is designed to preserve foundational representations while reducing computational overhead during fine-tuning. Beyond these, several other PEFT techniques have been applied to Code2Code tasks, where code manipulation is central. Prefix Tuning is adopted in 2 studies (Choi and Lee, 2023; Liu et al., 2023b), while Prompt Tuning is utilized in 3 studies (Liu et al., 2023b; Wang et al., 2023c, 2022b).
@@
LoRA-based Base LoRA APR(Li et al., 2024b; Liu et al., 2024c; Huang et al., 2025), CComp(Liu et al., 2024a), CG(Fakih et al., 2024; Hajipour et al., 2024; Ayupov and Chirkova, 2022; Wang et al., 2025; Yang et al., 2024a; Liu et al., 2024a), CMG(Fan et al., 2024; Liu et al., 2024b), CRF(Hajipour et al., 2024), CRG(Fan et al., 2024), CSum(Liu et al., 2023b; Ayupov and Chirkova, 2022), CTrans(Liu et al., 2023b; Ayupov and Chirkova, 2022), JITCU(Fan et al., 2024) CCD(Liu et al., 2023b; Ayupov and Chirkova, 2022), CR(Yu et al., 2024), DD(Liu et al., 2024b, 2023b)
@@
LoRA: Among all PEFT techniques, LoRA emerges as the most widely adopted approach across software engineering tasks, being applied in 35 distinct SE task instances. When including its hybrid variant, FF-LoRA combined with Adapter Tuning, the total number of applications extends to 39 SE task instances. To uncover how LoRA is typically configured, we present a synthesis of its use across both generative and non-generative SE tasks in Table 3. The table captures commonly used rank values, target layers (e.g., Attention, Feed-Forward), and the proportion of trainable parameters. By contrasting configurations across task categories, we identify trends that reflect LoRA’s adaptability and efficiency in varied contexts.
@@
Wang et al. (Wang et al., 2025) CG 8 16 $\times$ 3.86%
@@
As shown in Table 5, Prefix Tuning has been applied across diverse SE tasks with considerable variation in configuration. In generative tasks, notable is the work by Choi et al. (Choi and Lee, 2023) in which Prefix Tuning has been explored for different scenarios, including Code Generation, Code Summarization, and Code Translation. For each specific tasks, the prefix lengths were adjusted to encode the naturally-occurring semantic of the task itself.
@@
Our analysis reveals consistent trends in the use of model architectures and PEFT methods across software engineering tasks. Encoder-decoder architectures are the most commonly used (29 studies), followed by encoder-only (22) and decoder-only (18). Among base models, CodeT5 appears most frequently (19 studies), particularly in generation tasks, followed by GraphCodeBERT (10), CodeBERT and PLBART (9 each), and UniXcoder (8). Variants such as CodeT5-base (10 studies) and the CodeLlama series (7B, 13B, 34B) show increasing adoption, along with emerging large-scale models like StarCoder-15B, DeepSeek-Coder-33B, and LLaMA-2-70B. Base LoRA is the most widely used PEFT technique (24 studies), especially in code generation and program repair. Prompt-based strategies, such as Prompt Tuning and Prefix Tuning, are widely adopted—especially in tasks like Code Summarization and Code Translation. Our in-depth analysis of three PEFT methods highlights that LoRA is often preferred for its strong balance between performance and parameter efficiency, while Prompt and Prefix Tuning are valued for their flexible input-conditioning capabilities. Notably, some studies explore hybrid configurations, combining Prefix Tuning with LoRA, suggesting early experimentation with multi-strategy adaptation. Despite these advances, configuration details are not consistently reported across the literature, particularly for adapter-based approaches, limiting reproducibility and comparative analysis.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
As anticipated, our analysis focuses exclusively on studies that benchmark PEFT methods against full fine-tuning, enabling controlled evaluations grounded in clearly reported performance and efficiency metrics. This selection criterion led to the inclusion of 17 studies in our evaluation. At the same time, it highlights a notable gap in the literature: $\sim$37% of the reviewed works do not include any direct comparison with full fine-tuning model, making it difficult to assess the practical trade-offs of PEFT methods in those cases.
@@
In generative tasks, the application of PEFT methods demonstrates to produce results on par to full model’s parameters fine-tuning or surprisingly even better in both accuracy and efficiency. Based on the evidence from the 11 generative task entries in Table 6, ~75% report performance improvements ($\blacktriangle$), and over half (~54%) also report gains in training or memory efficiency ($\blacktriangle$). These trends are most pronounced in Code Summarization, Code Generation, and Code Refinement, which collectively represent the most frequently studied generative SE tasks in PEFT evaluations.
[…]
@@
On different, yet related note, our analysis of Table 6 reveals a notable trend in the literature: researchers are increasingly employing PEFT methods not only to improve fine-tuning efficiency but also to enhance model performance on the target task. While this may initially appear counterintuitive—given that PEFT techniques are typically associated with reducing computational overhead—our systematic review identified at least 7 studies across generative tasks where the primary motivation was performance optimization. In these studies, PEFT techniques such as LoRA, QLoRA, Prompt Tuning, and Adapter Fusion were adopted with the explicit goal of maximizing task-specific accuracy or generalization, rather than solely minimizing resource usage. This dual role of PEFT—as a strategy for achieving both computational efficiency and performance improvements, or the latter alone–communicates its versatility and growing relevance in generative software engineering tasks. As the demand for deploying large language models in real-world development workflows continues to rise, PEFT emerges as a practical and scalable alternative to full fine-tuning, offering significant gains in both accuracy and adaptability across diverse code generation scenarios.
@@
Our analysis reveals significant disparities in PEFT method coverage across software engineering tasks. While LoRA-based and adapter-based demonstrate consistent effectiveness (both in generative and non-generative tasks), we note a striking absence of empirical evaluations for several established PEFT methods—including AdaLoRA, Prefix Tuning, P-Tuning, and BitFit. The current evidence strongly favors LoRA variants for Code Generation, whereas non-generative tasks show more varied outcomes, with adapter-based emerging as the most used approach.
@@
While current evidence demonstrates the effectiveness of certain PEFT methods (particularly LoRA variants and Adapter-based) for software engineering tasks, the limited evaluation of other promising approaches like AdaLoRA, Prefix-Tuning, P-Tuning, and BitFit leaves their potential largely unexplored. Future work should expand the methodological diversity of PEFT evaluations through standardized benchmarking across different task categories, which would enable more comprehensive comparisons and help identify optimal adaptation strategies for specific SE scenarios. Such efforts could particularly benefit from investigating how understudied methods might address current limitations in parameter efficiency and task performance, while establishing clearer guidelines for method selection based on task characteristics and resource constraints.

### v3

@@ 4.2.2. PEFT Methods Across Software Engineering Tasks and Model Architectures
In this section, we explore how PEFT methods are distributed across various software engineering tasks, distinguishing them between generative and non-generative. Among the identified techniques, Base LoRA is the most dominant, appearing in 29 SE tasks. The Adapter family has been applied across 23 software engineering tasks, encompassing specialized variants such as L-Adapters and T-Adapters. Other frequently used techniques include Prompt Tuning (14 studies), Prefix Tuning (13), and P-Tuning (7), which leverage soft or continuous prompts to inject task-specific knowledge into frozen pre-trained models. Collectively, these methods demonstrate significant effectiveness in reducing training costs while maintaining or enhancing performance, affirming their value in scalable and efficient model adaptation across a diverse set of software engineering tasks.
@@
Code Generation is addressed in 10 studies, making it one of the most frequently explored generative SE tasks in the context of PEFT. LoRA remains the dominant technique, appearing in 7 studies (Wang et al., 2025b; Liu et al., 2024a; Fakih et al., 2024; Hajipour et al., 2024; Ayupov and Chirkova, 2022; Yang et al., 2024a; Wang et al., 2025a), including combinations of FF-LoRA variants with Adapter Tuning, one study implemented Code Summarization, Code Clone Detection, Code Generation, and Code Translation (Ayupov and Chirkova, 2022).
@@
A further application of SE automation revolves around the generation of unit tests–referred to as Unit Test Generation in our taxonomy Fig. 3. In particular, for such a task we have found two distinct studies, each employing distinct PEFT approaches. Liu et al. (Liu et al., 2024a) applied QLoRA within a multi-task fine-tuning framework (MFT-5Tasks), while Zlotchevski et al. (Zlotchevski et al., 2022) employed Prefix Tuning, wherein 200 trainable virtual tokens were prepended to a frozen BART model.
@@
Looking at SE tasks for which code changes optimization is requested, we find a surprisingly widespread application of PEFT techniques to support Code Refinement activities (i.e., bug-fixing). This peculiar task, has been explored through a diverse range of PEFT techniques, which—consistent with findings from Code Generation—span the full spectrum of parameter-efficient optimization strategies. In this context, Nuo et al. (Chen et al., 2023b) introduced Pass-Tuning, a lightweight method that tunes less than 1% of parameters by leveraging U-AST (Upgraded Abstract Syntax Tree) token sequences within a graph-attentional model. Building on structure-aware prompting, Yun et al. (Peng et al., 2024) applied Prompt Tuning to CodeT5-base, crafting domain-specific prompts based on AST-derived fix templates. Saberi et al. (Saberi and Fard, 2023) proposed a fusion-based approach, combining a Java-specific L-Adapter with a token-type-aware NER-Adapter via AdapterFusion to improve bug-fixing effectiveness. Similarly, Hajipour et al. (Hajipour et al., 2024) investigated the use of LoRA as a parameter-efficient alternative for Code Refinement tasks. Traditional Prompt Tuning appears in one study (Peng et al., 2024), while a multi-perspective adaptation strategy—incorporating L-Adapter, NER-Adapter, and AdapterFusion—is explored in a different research (Saberi and Fard, 2023). Additionally, one study applies a comprehensive suite of PEFT methods—including Pass-Tuning, BitFit, Adapters, and P-Tuning—to enhance source code refinement (Chen et al., 2023b).
@@
Still in the domain of Code2Code tasks, Code Translation is addressed in 8 studies (Choi and Lee, 2023; Chen et al., 2023b; Liu et al., 2023b; Wang et al., 2023c; Ayupov and Chirkova, 2022; Wang et al., 2022b; Liu et al., 2024a; Wang et al., 2025a), while Code Completion is explored in 3 studies (Shi et al., 2023; Liu et al., 2024a; Wang et al., 2025a). Notably, for the latter, Shi et al. (Shi et al., 2023) introduce a customized PEFT technique called Telly-K, which modifies the training procedure by selectively freezing the lower layers of the model. This approach is designed to preserve foundational representations while reducing computational overhead during fine-tuning. Beyond these, several other PEFT techniques have been applied to Code2Code tasks, where code manipulation is central. Prefix Tuning is adopted in 2 studies (Choi and Lee, 2023; Liu et al., 2023b), while Prompt Tuning is utilized in 3 studies (Liu et al., 2023b; Wang et al., 2023c, 2022b).
@@
LoRA-based Base LoRA APR(Li et al., 2024b; Liu et al., 2024c; Huang et al., 2025), CComp(Liu et al., 2024a) (Wang et al., 2025a), CG(Fakih et al., 2024; Hajipour et al., 2024; Ayupov and Chirkova, 2022; Wang et al., 2025b; Yang et al., 2024a; Liu et al., 2024a), (Wang et al., 2025a) CMG(Fan et al., 2024; Liu et al., 2024b), CRF(Hajipour et al., 2024), CRG(Fan et al., 2024), CSum(Liu et al., 2023b; Ayupov and Chirkova, 2022), CTrans(Liu et al., 2023b; Ayupov and Chirkova, 2022) (Wang et al., 2025a), JITCU(Fan et al., 2024), PBT(Wang et al., 2025a) CCD(Liu et al., 2023b; Ayupov and Chirkova, 2022), CR(Yu et al., 2024), DD(Liu et al., 2024b, 2023b), HFP(Wang et al., 2025a)
@@
LoRA: Among all PEFT techniques, LoRA emerges as the most widely adopted approach across software engineering tasks, being applied in 36 distinct SE task instances. When including its hybrid variant, FF-LoRA combined with Adapter Tuning, the total number of applications extends to 40 SE task instances. To uncover how LoRA is typically configured, we present a synthesis of its use across both generative and non-generative SE tasks in Table 3. The table captures commonly used rank values, target layers (e.g., Attention, Feed-Forward), and the proportion of trainable parameters. By contrasting configurations across task categories, we identify trends that reflect LoRA’s adaptability and efficiency in varied contexts.
@@
Wang et al. (Wang et al., 2025b) CG 8 16 $\times$ 3.86%
@@
As shown in Table 5, Prefix Tuning has been applied across diverse SE tasks with considerable variation in configuration. In generative tasks, notable is the work by Choi et al. (Choi and Lee, 2023) in which Prefix Tuning has been explored for different scenarios, including Code Generation, Code Summarization, and Code Translation. For each specific task, the prefix lengths were adjusted to encode the naturally-occurring semantics of the task itself.
@@
Our analysis reveals consistent trends in the use of model architectures and PEFT methods across software engineering tasks. Encoder-decoder architectures are the most commonly used (29 studies), followed by encoder-only (22) and decoder-only (19). Among base models, CodeT5 appears most frequently (19 studies), particularly in generation tasks, followed by GraphCodeBERT (10), CodeBERT and PLBART (9 each), and UniXcoder (8). Variants such as CodeT5-base (10 studies) and the CodeLlama series (7B, 13B, 34B) show increasing adoption, along with emerging large-scale models like StarCoder-15B, DeepSeek-Coder-33B, and LLaMA-2-70B. Base LoRA is the most widely used PEFT technique (29 studies), especially in code generation and program repair. Prompt-based strategies, such as Prompt Tuning and Prefix Tuning, are widely adopted—especially in tasks like Code Summarization and Code Translation. Our in-depth analysis of three PEFT methods highlights that LoRA is often preferred for its strong balance between performance and parameter efficiency, while Prompt and Prefix Tuning are valued for their flexible input-conditioning capabilities. Notably, some studies explore hybrid configurations, combining Prefix Tuning with LoRA, suggesting early experimentation with multi-strategy adaptation. Despite these advances, configuration details are not consistently reported across the literature, particularly for adapter-based approaches, limiting reproducibility and comparative analysis. Architecture-task alignment analysis reveals that decoder-only models appeared exclusively in generative tasks, suggesting fundamental architectural requirements for different SE task categories and indicating a strong correlation between model design and task functionality.
@@
Table 6. Summary of Best-Performing PEFT Methods Across Software Engineering Tasks. We report both performance impact ( P= improved, P= declined, P= similar) and training/memory efficiency ( E= more efficient, E= less efficient, E= similar), based on comparisons with corresponding full fine-tuning baselines. If only performance is reported (e.g., P), it indicates that efficiency was not addressed in the original study. Each entry represents the best-performing PEFT variant identified in the respective study for the given task and base model. Tasks are grouped into generative and non-generative categories for clarity.
@@
LoRA-based Base LoRA APR(Huang et al., 2025) P ECG(Wang et al., 2025b) P E, CG(Fakih et al., 2024) P E, CG(Hajipour et al., 2024) P, CG(Yang et al., 2024a) P, CG(Wang et al., 2025a) P, CComp(Wang et al., 2025a) P, CRF(Hajipour et al., 2024) P, CSum(Ayupov and Chirkova, 2022) P, CTrans(Wang et al., 2025a) P, PBT(Wang et al., 2025a) P CCD(Ayupov and Chirkova, 2022) P, CR(Yu et al., 2024) P, DD(Liu et al., 2024b) P, HFP(Wang et al., 2025a) P
QLoRA CComp(Liu et al., 2024a) P E, CG(Liu et al., 2024a) P E, CSum(Afrin et al., 2025) P E $\times$
@@
Adapter-based Base Adapter CMG(Liu et al., 2024b) P E, CSum(Wang et al., 2023a) P, CTrans(Chen et al., 2023b) P CCD(Chen et al., 2023b) P, CCD(Liu et al., 2023b) P, CS(Wang et al., 2023a) P
L-Adapter $\times$ CT(Saberi et al., 2024) P E
T-Adapter CSum(Saberi et al., 2024) P E $\times$
Prompt-based Prompt Tuning CSum(Wang et al., 2023c) P, CSum(Wang et al., 2022b) P, CTrans(Wang et al., 2023c) P, CTrans(Wang et al., 2022b) P CS(Wang et al., 2023c) P, DD(Wang et al., 2023c) P, DD(Wang et al., 2022b) P
@@
Pass-Tuning CG(Chen et al., 2023b) P, CRF(Chen et al., 2023b) P, CSum(Chen et al., 2023b) P $\times$
(IA)3 APR(Li et al., 2024b) P E $\times$
@@
FF-LoRA + AT CTrans(Ayupov and Chirkova, 2022) P CCD(Ayupov and Chirkova, 2022) P
L-Adapter + T-Adapter $\times$ CCD(Saberi et al., 2024) P E
L-Adapter + NER-Adapter + AdapterFusion CSum(Saberi and Fard, 2023) P E, CRF(Saberi and Fard, 2023) P E $\times$
Telly-K CComp(Shi et al., 2023) P E, CG(Shi et al., 2023) P E, CSum(Shi et al., 2023) P E CCD(Shi et al., 2023) P E, CS(Shi et al., 2023) P E
MAM CSum(Liu et al., 2023b) P, CTrans(Liu et al., 2023b) P DD(Liu et al., 2023b) P
@@
As anticipated, our analysis focuses exclusively on studies that benchmark PEFT methods against full fine-tuning, enabling controlled evaluations grounded in clearly reported performance and efficiency metrics. This selection criterion led to the inclusion of 20 studies in our evaluation. At the same time, it highlights a notable gap in the literature: $\sim$29% of the reviewed works do not include any direct comparison with full fine-tuning model, making it difficult to assess the practical trade-offs of PEFT methods in those cases.
@@
In generative tasks, the application of PEFT methods demonstrates to produce results on par to full model’s parameters fine-tuning or surprisingly even better in both accuracy and efficiency. Based on the evidence from the 34 generative task entries in Table 6, ~72% report performance improvements ( P), and over half (~54%) also report gains in training or memory efficiency ( E). These trends are most pronounced in Code Summarization, Code Generation, and Code Refinement, which collectively represent the most frequently studied generative SE tasks in PEFT evaluations.
Among the various strategies, LoRA-based methods—particularly Base LoRA and QLoRA—stand out as the most widely adopted and effective. Base LoRA is applied across four distinct generative tasks: i.e., Automated Program Repair, Code Generation, Code Refinement, and Code Summarization. All cases report improved task performance ( P), with efficiency gains ( E) observed in APR and one Code Generation study, while one Code Generation study reports a slight drop in efficiency ( E). Similarly, QLoRA shows high effectiveness in Code Generation, Code Completion, and Code Summarization, consistently yielding performance and efficiency gains ( P E), further solidifying the role of LoRA-based methods as leading choices for token-level generation tasks. Recent comprehensive evaluations by Wang et al. (Wang et al., 2025a) provide additional empirical validation of these performance patterns while revealing important limitations in domain-specific scenarios. Their systematic comparison across five code intelligence tasks demonstrates that while PEFT methods like LoRA can achieve competitive performance on public datasets, substantial performance gaps emerge when applied to private data contexts. Their findings reveal mixed effectiveness of PEFT methods, with competitive performance on general-domain tasks but notable performance degradation in scenarios involving private codebases and specialized domain knowledge, where the distribution shift from pre-training data was substantial. This evidence reinforces our systematic findings that while LoRA demonstrates overall effectiveness across diverse applications, significant challenges remain in complex scenarios where substantial domain adaptation and knowledge acquisition beyond the model’s pre-training data are required. Adapter-based methods show more varied results. In particular, Base Adapter achieves performance improvements in Code Summarization and Commit Message Generation—with the latter performing on par to the full fine-tuned model when evaluated for generating code summaries ( P), but gaining in efficiency ( E). However, its application to Code Translation results in a performance decline ( P), a trend mirrored by the hybrid FF-LoRA + AT and MAM approaches. These declines across three different studies indicate that Code Translation may be less amenable to parameter-efficient strategies, likely due to its higher sensitivity to sequence-level alignment and cross-language mapping.
On the other hand, Prompt-based methods yield generally strong results in generative contexts. Prompt Tuning is successfully applied to Code Summarization and Code Translation, achieving consistent performance gains ( P). Likewise, Pass-Tuning and (IA)3 also yield improvements in specific tasks—Code Refinement, Code Summarization, and Automated Program Repair, respectively—with (IA)3 additionally reporting efficiency gains ( E). However, Pass-Tuning shows a performance decline ( P) in one Code Generation instance, suggesting that tuning strategies may not be uniformly effective across all task types or data distributions.
Hybrid and other PEFT methods such as Telly-K, L-Adapter + NER-Adapter + AdapterFusion demonstrate promising performance, particularly in Code Summarization and Code Refinement. Notably, Telly-K reports performance and efficiency gains ( P E) in three generative tasks—Code Generation, Code Summarization, and Code Completion—suggesting that selective fine-tuning of higher transformer layers can effectively retain model expressiveness while reducing overhead. In contrast, MAM shows performance degradation ( P) in both Code Summarization and Code Translation. This may be attributed to limitations in modular adaptation when the task requires transforming data from one domain and projecting it onto a different target domain (e.g., translate code into natural language).
In non-generative tasks, PEFT methods demonstrate a more nuanced performance profile. Based on the 16 non-generative task entries in Table 6, approximately $\sim$62% report improved task performance ( P), while around $\sim$50% show gains in efficiency ( E). This reflects moderate but meaningful benefits of PEFT in retrieval and classification SE tasks. The most frequent task is Code Clone Detection, which shows mixed results. In this case, we count three entries where improved or neutral performance, three show performance degradation ( P), indicating sensitivity to semantic precision and token similarity under compressed training regimes. In contrast, Defect Detection tasks demonstrate consistent improvements in performance, particularly when Prompt Tuning or LoRA is used.
Among PEFT strategies, LoRA-based methods (e.g., Base LoRA) report success in Defect Detection and Code Review, with consistent performance improvements ( P). However, its application to Code Clone Detection shows performance degradation ( P), suggesting potential limitations in handling fine-grained syntactic variance. Adapter-based methods yield mixed outcomes. For instance, Base Adapter leads to performance improvements in Code Search, but its results in Code Clone Detection are not so evident—improving in one study and degrading in another. The application of L-Adapter to the Cloze Test task yields improvements in both performance and efficiency ( P E), indicating a promising outcome for structured classification tasks. Prompt-based methods, especially Prompt Tuning, yield consistent performance gains in Defect Detection and Code Search, aligning well with their token-efficient query-focused design. Notably, these methods are not applied to Code Clone Detection or Code Review in the current set of studies, leaving room for future exploration. Finally, hybrid and selective tuning methods, such as Telly-K and L-Adapter + T-Adapter, demonstrate performance gains in tasks like Code Clone Detection and Code Search ( P E).
@@
On a different, yet related note, our analysis of Table 6 reveals a notable trend in the literature: researchers are increasingly employing PEFT methods not only to improve fine-tuning efficiency but also to enhance model performance on the target task. While this may initially appear counterintuitive—given that PEFT techniques are typically associated with reducing computational overhead—our systematic review identified at least 7 studies across generative tasks where the primary motivation was performance optimization. In these studies, PEFT techniques such as LoRA, QLoRA, Prompt Tuning, and Adapter Fusion were adopted with the explicit goal of maximizing task-specific accuracy or generalization, rather than solely minimizing resource usage. This dual role of PEFT—as a strategy for achieving both computational efficiency and performance improvements, or the latter alone–communicates its versatility and growing relevance in generative software engineering tasks. On an interesting note, the recent paper by Wang et al. (Wang et al., 2025a) provides valuable insights into the effectiveness of PEFT methods for software engineering across both private (non–open-source) and public (open-source) application domains. Their findings show that, while PEFT methods perform competitively on general-domain tasks, performance declines markedly in scenarios requiring substantial adaptation to previously unseen knowledge. This suggests that PEFT’s effectiveness may be inherently constrained by the degree of domain-specific adaptation needed. However, this outcome may also be influenced by the predominant reliance of large code models on open-source training data–typically sourced from public GitHub repositories. As a result, the observed performance drop could partly stem from a data distribution shift caused by applying PEFT across two distinct distributions: the pre-training dataset and the domain-specific data used for PEFT-optimized fine-tuning. As the demand for deploying large language models in real-world development workflows continues to rise, PEFT emerges as a practical and scalable alternative to full fine-tuning, offering significant gains in both accuracy and adaptability across diverse code generation scenarios.
@@
Our analysis reveals significant disparities in PEFT method coverage across software engineering tasks. While LoRA-based and adapter-based demonstrate consistent effectiveness (both in generative and non-generative tasks), we note a striking absence of empirical evaluations for several established PEFT methods—including AdaLoRA, Prefix Tuning, P-Tuning, and BitFit. The current evidence strongly favors LoRA variants for Code Generation, whereas non-generative tasks show more varied outcomes, with adapter-based emerging as the most used approach. Performance analysis reveals that $\sim$72.7% of generative tasks report performance improvements compared to $\sim$62.5% of non-generative tasks, demonstrating fundamental differences in PEFT applicability across task categories. Efficiency-performance correlation analysis shows dual benefits, with over half of generative tasks achieving efficiency gains alongside performance improvements, while around $\sim$25% of non-generative tasks demonstrate efficiency gains. Critical analysis of Base LoRA shows $\sim$60% success rates against full fine-tuning baselines, challenging conventional assumptions about parameter efficiency trade-offs and suggesting that targeted parameter updates may be more effective than comprehensive fine-tuning for specific software engineering contexts.
@@
While current evidence demonstrates the effectiveness of certain PEFT methods (particularly LoRA variants and Adapter-based) for software engineering tasks, the limited evaluation of other promising approaches like AdaLoRA, Prefix-Tuning, P-Tuning, and BitFit leaves their potential largely unexplored. Future work should expand the methodological diversity of PEFT evaluations through standardized benchmarking across different task categories, which would enable more comprehensive comparisons and help identify optimal adaptation strategies for specific SE scenarios. Such efforts could particularly benefit from investigating how understudied methods might address current limitations in parameter efficiency and task performance, while establishing clearer guidelines for method selection based on task characteristics and resource constraints. Additionally, future research must address domain-specific PEFT optimization that accounts for software engineering’s unique characteristics. This enhancement should include recommendations for adapting PEFT methods to preserve code structure and formatting requirements, providing evidence-based guidance that distinguishes between methods suitable for structured code generation versus natural language synthesis in software engineering contexts. Research should emphasize the importance of code-specific evaluation metrics such as syntactic correctness and compilability alongside traditional performance measures, while investigating PEFT methods specifically designed for code’s hierarchical and syntactic properties. This comprehensive approach will provide the domain-aware guidance necessary for effective PEFT adaptation in software engineering, recognizing that methods optimized for general natural language processing may require significant modification to handle the structured, syntax-dependent nature of code and software engineering artifacts.
5. Practical Implications and Strategic Recommendations
Building on our systematic analysis of PEFT applications in software engineering, we synthesize our findings into actionable insights across several critical dimensions: methodological strategy, resource optimization.
Methodological Strategy Implications: The documented concentration in synthesis-based applications suggests that organizations should prioritize PEFT investment in code generation and summarization infrastructure, as these represent the most mature and validated use cases. The exclusive association of decoder-only architectures with generative tasks provides clear architectural selection criteria for practitioners. Organizations implementing code generation systems should prioritize decoder-only models, while those requiring mixed analytical and generative capabilities should consider encoder-decoder architectures. The demonstrated reliability of Base LoRA across diverse applications establishes it as a primary choice for initial PEFT implementations, particularly when performance parity with full fine-tuning is critical.
Resource Optimization Implications: The superior performance-efficiency correlation in generative tasks suggests that organizations should prioritize PEFT adoption for generative applications to maximize return on computational investment. The lower but still substantial benefits in non-generative tasks indicate that PEFT adoption should follow a staged approach, beginning with generative applications before expanding to analytical tasks. The evidence for performance enhancement beyond traditional fine-tuning suggests that PEFT should not be viewed merely as a resource-saving compromise, but as a potentially superior adaptation strategy. This paradigm shift has significant implications for model development budgets and computational resource allocation in software engineering organizations. The current state of PEFT research demonstrates remarkable community progress, with widespread availability of optimized implementations through platforms like HuggingFace PEFT, facilitating practical adoption and reproducibility across software engineering applications.
Our findings demonstrate that PEFT optimization methods, including both parameter-efficient approaches like LoRA and prompt-based methods like Prompt Tuning and Prefix Tuning, often match or exceed full fine-tuning performance while requiring significantly fewer computational resources. This enables researchers to establish strong performance baselines through PEFT evaluation without requiring computationally expensive full model fine-tuning, promoting both research efficiency and environmental sustainability in large language model adaptation for software engineering tasks.